OpenStudy (anonymous):

The equation shows one mole of ethanol fuel being burned in oxygen. Convert the energy released into its equivalent mass. C2H5OH(l) + 3 O2(g) Æ 2 CO2(g) + 3 H2O (l) ΔH = -1418 kJ/mol

OpenStudy (anonymous):

deltaH=E E=m*c^2 --> m= E/(c^2) = 1418 *10^3 / (3*10^8)^2